# label=0.,leftmargin=15pt,labelwidth=10pt,labelsep=5pt, topsep=0pt,parsep=0pt,partopsep=0pt,noitemsep

### v1

@@ 1 Introduction
The initial work on V:N:M sparsity primarily focuses on designing its acceleration kernels, with only basic verification of its effect on the accuracy of BERT using the SQuAD task. However, the impact of V:N:M sparsity on broader Transformer models, such as vision Transformers and large language models (LLMs) like the Llama series, remains under-explored. Additionally, fundamental issues regarding V:N:M sparsity, such as how to select appropriate values for V and M in a Transformer architecture, remain unresolved. In this work, we aim to bridge these gaps by systematically investigating the application of V:N:M sparsity across multiple Transformer models, with a particular emphasis on enhancing their accuracy. Specifically, our contributions are as follows:
  • We propose a scheme that enables the generation of highly accurate V:N:M-sparse Transformers under different constraints, which broadens the applicability of V:N:M sparsity.
@@ 3 Preliminary
In Eq. 2 and 3, $\nabla\mathcal{L}$ denotes the gradient, while $\gamma$ and $\lambda$ are the learning rate and regularization coefficient, respectively. Eq. 2 indicates that only the retained weights are updated, with the V:N:M-sparse mask $\bm{\mathbf{M}}$ remaining unchanged after one-shot pruning. In contrast, Eq. 3 gradient-updates the dense $\bm{\mathbf{W}}$ and $\bm{\mathbf{M}}$ is time-variant.
@@ 4 Methods
Scheme overview We show the process of generating a V:N:M-sparse Transformer with high accuracy in Figure 2(a). Given a pretrained dense Transformer and a specified speedup threshold, a heuristic approach is employed to select appropriate V and M values for pruning the dense Transformer. After that, we consider two distinct scenarios. In the first scenario, where a limited training budget is available, V:N:M-specific CP, RIA-based pruning, and fixed mask training are sequentially employed. CP and RIA can significantly improve the accuracy of V:N:M-sparse Transformers upon pruning, while for low training budget constraints, fixed mask training, no matter with full-parameters or LoRA, incurs significantly lower overhead compared to dynamic mask training as the mask update costs are canceled. Note the necessity of RIA is further explained in Appendix B.
@@ 4.1 V and M selection
It is straightforward to prove that for the same Transformer, the relative order of different $MD_{f}$ is entirely determined by the values of V and M, i.e., $K$ in Eq. 5, and is irrespective of the weight shapes of the linear layers (See Appendix C for the proof). Thus, only calculating $K$ suffices for comparison and choosing the best (V, M) combination.
@@ 4.3 Three-staged LoRA training
where $\bm{\mathbf{B}}$ and $\bm{\mathbf{A}}$ are two low rank matrices. Normally, $Mv$ is a function of $Wv$, $Bv$, and $Av$. During training, $Wv$ remains fixed while $Bv$ and $Av$ are updated. We propose a three-stage LoRA training technique to enable dynamic mask training with LoRA and enhance the accuracy of V:N:M-sparse Transformers. 1) Dense LoRA. At the beginning of training, standard LoRA fine-tuning is applied to the Transformer, where the masks $\bm{\mathbf{M}}$ are consistently all-one matrices. 2)
@@
Sparse LoRA with the dynamic masks. After the dense LoRA, the masks $\bm{\mathbf{M}}$ are updated at regular intervals according to the V:N:M sparse patterns as training progresses, which means each update occurs after a fixed number of iterations. During the mask update, the low-rank matrices $\bm{\mathbf{B}}$ and $\bm{\mathbf{A}}$ are merged with the original dense weight matrix $\bm{\mathbf{W}}$ before calculating the importance scores and conducting V:N:M-sparse pruning. 3) Sparse LoRA with the fixed masks. While dynamic mask updates facilitate the exploration of appropriate V:N:M-sparse masks, they can also introduce instability in the training process. Furthermore, directly applying regularization such as SR-STE, as illustrated in Eq. 3, is challenging because the LoRA term $\bm{\mathbf{B}}\bm{\mathbf{A}}$ is complex to regularize. Therefore, in the third stage, we advocate for fine-tuning with fixed masks to balance the exploration and exploitation of masks. At this stage, the masks $\bm{\mathbf{M}}$ are inherited from the last update in the previous stage and remain unchanged until the training is completed.
It is important to note that the number of iterations in the first two stages should only constitute a smaller proportion of the total iterations, with the majority allocated to the third stage, i.e., Sparse LoRA with fixed masks. This is because, in the absence of regularization, frequent updates to the masks can negatively impact the gradient flow of V:N:M sparse Transformers during fine-tuning, as exemplified in Figure 3. In practice, the iterations for the first two stages should not exceed 10% of the total iterations.
@@ 5 Experiments
Models, datasets and tasks To evaluate the proposed V:N:M-sparse Transformer generation method—which incorporates V and M selection, V:N:M-specific channel permutations, and three staged LoRA training techniques, three benchmarks have been established: 1) DeiT (Touvron et al., 2021) for image classification. This benchmark is widely recognized for assessing the efficacy of model compression techniques in vision Transformers. Given that V:N:M sparsity operates at high sparsity levels (greater than 50%), the DeiT-tiny is excluded due to insufficient redundancy. The datasets used for the tasks include ImageNet-1K (Deng et al., 2009), Cifar-10 and Cifar-100 Krizhevsky et al. (2009), Bird and Vehicle from the subset of ImageNet-1K. Note that the latter four datasets are used to form downstream tasks. 2) Swin Transformers. This category of vision Transformers, known for its hierarchical architecture and shifted window mechanisms, demonstrates increased sensitivity to model compression (Liu et al., 2021). In this work, the V:N:M-sparse Swin Transformers are assessed across two tasks: image classification on the ImageNet-1K dataset and object detection on the COCO 2017 dataset (Lin et al., 2014). 3) Llama2-7B on downstream tasks including predicting the next token on wikitext2 (Merity et al., 2016) and eight well-established 5-shot tasks (Gao et al., 2021). Note that Llama2-7B undergoes LoRA training on both Wikitext2 and Alpaca (Taori et al., 2023) before testing on these tasks, which is a standard practice in the fine-tuning of LLMs. In addition, the speedups of these V:N:M-sparse Transformers compared to their dense counterparts are measured on RTX 3090 GPUs, which were also the speed-testing platform in the initial study.
@@ 5.1 Results for V and M selection
Figure 4 compares the accuracy and speedup for V:N:M-sparse Transformers, both with and without the proposed V and M selection technique. In this experiment, for practical speedups, V is confined to the values $[16,32,64,128]$. Sparse Transformers with varying M values and unified V=64 are selected to establish the speedup threshold $s$, as specified in Eq. 4, with 64 representing a central value in the V distribution. Using the thresholds, the V and M selection technique is applied to derive new (V, M) and further generate the V:N:M-sparse Transformers accordingly, as indicated by the yellow lines in Figure 4.
Specifically, in Figure 4(a), the V:N:M-sparse DeiT-base models located on the Pareto front exhibit higher Top-1 accuracy compared to the baseline, achieving a maximum accuracy improvement of 2.6% under similar speedup conditions. For the Llama2-7B, shown in Figure 4(b), the maximal average score difference is 3.41. Besides, when appropriate values of V and M are utilized, the sparse DeiT-base maintains a nearly lossless accuracy of 81.59% while achieving a speedup of 2.02 relative to its dense counterpart. Similarly, the sparse Llama2-7B achieves a speedup of 1.65, with a score of 50.85 on 5-shot tasks. Furthermore, it is essential to emphasize that V:N:M-sparse vision Transformers exhibit substantially greater speedups compared to 2:4 counterparts. As depicted in Figure 4(a), the 128:2:6-sparse DeiT-base matches the accuracy of the 2:4-sparse DeiT-base while achieving a 1.71x speedup over its dense counterpart, in contrast to the 2:4-sparse version, which achieves only a 1.15x speedup. More speedup results of V:N:M-sparse Transformers are shown in Appendix E.
Figure 4: The speedup-accuracy curves of V:N:M-sparse Transformer. a) Top-1 accuracy of DeiT-base using TS2 with different V and M values. b) Average scores of Llama2-7B using TS3 on 5-shot tasks with different V and M values. The speedup-accuracy curves using TS1 are shown in Figure 7 in Appendix D.
@@ 5.2 Results for V:N:M-specific channel permutation and TS1
Since the proposed V:N:M-specific CP is effective for low-budget training scenarios, we assess this technique on downstream tasks that allow for a limited number of training epochs to attain acceptable accuracy. The results, presented in Tables 3 and 4, were all obtained using TS1. For vision downstream tasks with a maximum of 30 training epochs, our technique enhances final accuracy by 0.71% for full-parameter training and 1.43% for LoRA training. Notably, when V:N:M-sparse pruning is only applied without any training, the accuracy gap can increase to 5.87% with or without our technique. Similarly, for Llama2-7B, which underwent only 1500 training iterations—a relatively short duration compared to the 50,000 iterations shown in the subsequent Table 8, our technique improves the 5-shot scores by 1.25.
@@ 5.3 Results for TS2
Table 5: Results comparison of sparse DeiTs on ImageNet-1K
Model Param. (M) (↓%) FLOPs (G) (↓%) Top-1 Acc. (%) $\Delta$
DeiT-B 86.6 - 17.6 - 81.84 -
DeiT-B-600e 86.6 - 17.6 - 82.01 +0.17
T2T-ViT-24 64.1 26.0 13.8 21.6 82.30 +0.46
PVT-L 61.4 29.1 9.8 44.3 81.70 -0.14
AutoFormer-B 54.0 37.6 11.0 37.5 82.40 +0.56
SSP-B 56.8 34.4 11.8 33.1 80.80 -1.04
S2ViTE-B 56.8 34.4 11.8 33.1 82.22 +0.38
Evo-ViT 56.3 - 11.7 33.3 80.30 -0.54
EViT-DeiT-B 86.6 - 11.5 34.7 81.50 -0.34
DynamicViT 61.0 - 11.5 34.7 81.30 -0.54
ViT-Slim 52.6 39.3 10.6 39.6 82.40 +0.56
VTP-B 47.3 45.4 10.0 43.2 80.70 -1.14
PS-ViT-B 86.6 - 9.8 44.3 81.50 -0.34
UVC - - 8.0 54.5 80.57 -1.27
SAViT 25.4 70.7 5.3 69.9 81.66 -0.18
Ours (64:2:8) 22.7 73.8 5.0 71.6 81.08 -0.76
Ours-600e (64:2:8) 22.7 73.8 5.0 71.6 81.76 -0.08
DeiT-S 22.1 - 4.6 - 79.85 -
DeiT-S-600e 22.1 - 4.6 - 80.02 +0.17
AdaViT-S 22.1 0.0 3.6 21.7 78.60 -1.25
DynamicViT 22.1 - 3.4 26.1 79.60 -0.25
EViT-DeiT-S 22.1 - 3.4 34.8 79.50 -0.35
SSP-S 14.6 33.3 3.1 31.6 77.74 -2.11
S2ViTE-S 14.6 33.3 3.1 31.6 79.22 -0.63
SAViT 14.7 33.5 3.1 31.7 80.11 +0.26
Ours (64:2:5) 9.3 57.9 2.1 54.3 78.97 -0.68
Ours-600e (64:2:5) 9.3 57.9 2.1 54.3 79.65 -0.2
@@
Table 6: V:N:M-sparse Swin Transformers on image classification and object detection
Model Method Param.(M) (↓%) FLOPs (G) (↓%) Top-1 Acuu. (%)
Swin-base Dense 87.8 - 15.4 - 83.51
LPViT 64.1* 27.0 11.24 27.0 81.7
Ours(32:2:5) 35.9 59.1 6.1 60.4 81.7
H-DETR (Swin-Tiny) Dense 40.2 - 212.0 - 45.3
Dense-24e 40.2 - 212.0 - 49.2
FMT(32:2:5) 18.1 55.0 93.7 55.8 47.5
Ours(32:2:5) 18.1 55.0 93.7 55.8 47.8
FMT(32:2:6) 15.5 61.4 78.0 63.2 44.1
Ours(32:2:6) 15.5 61.4 78.0 63.2 45.3
@@ 5.4 Results for three-staged LoRA training and TS3
Figure 5: Average scores of V:N:M-sparse Llama2-7B on 5-shot tasks under different LoRA finetuning schemes
@@ 6 Conclusion
This study focuses on enhancing the accuracy and accuracy-speedup trade-offs of V:N:M-sparse Transformers in multiple scenarios. We address the crucial yet unexplored questions specific to V:N:M sparsity, including selecting appropriate values for V and M, and CP tailored for V:N:M sparsity. Additionally, we propose a three-staged LoRA training technique, which for the first extends V:N:M sparsity to LLMs. Extensive experiments demonstrate that, with our methodology, V:N:M-sparse Transformers can attain nearly lossless accuracy or perform comparably to those with post-pruning 2:4 sparsity. Given its superior speed performance, we conclude that V:N:M sparsity is more effective than 2:4 for compressing highly redundant Transformers in inference-cost-sensitive scenarios. We hope our work will promote the widespread use of V:N:M sparsity as a truly effective solution for compressing Transformers.

### v2

@@ 1 Introduction
The initial work on V:N:M sparsity primarily focuses on designing its acceleration kernel. Importantly, the impact of V:N:M sparsity on broader Transformer models, such as vision Transformers and large language models (LLMs) like the Llama series, remains under-explored. Additionally, fundamental issues regarding V:N:M sparsity, such as how to select appropriate values for V and M in a Transformer architecture, have never been resolved. Without addressing these issues, V:N:M sparsity can not comprehensively outperform 2:4 sparsity in compressing Transformers with high redundancy. In this work, we aim to bridge these gaps by systematically investigating the application of V:N:M sparsity across multiple Transformer models, with a particular emphasis on enhancing their accuracy. Specifically, our contributions are as follows:
  • We propose a framework to enable the generation of highly accurate V:N:M-sparse Transformers under different constraints, which broadens the applicability of V:N:M sparsity.
@@ 3 Preliminary
In Eq. 2 and 3, $\nabla\mathcal{L}$ denotes the gradient, while $\gamma$ and $\lambda$ are the learning rate and regularization coefficient, respectively. $\overline{\bm{\mathbf{M}}}_{t-1}$ represents the logical not operation of $\bm{\mathbf{M}}_{t-1}$ at $t-1$, which enables regularization to only target the pruned weights and gradually decrease their norms. Eq. 2 indicates that only the retained weights are updated, with the V:N:M-sparse mask $\bm{\mathbf{M}}$ remaining unchanged after one-shot pruning. In contrast, Eq. 3 gradient-updates the dense $\bm{\mathbf{W}}$ and $\bm{\mathbf{M}}$ is time-variant.
@@ 4 Methods
Scheme overview We show the process of generating a V:N:M-sparse Transformer with high accuracy in Figure 2(a). Given a pretrained dense Transformer and a specified speedup threshold, a heuristic approach is employed to select appropriate V and M values for pruning the dense Transformer. After that, we consider two distinct scenarios. In the first scenario, where a limited training budget is available, V:N:M-specific CP, RIA-based pruning, and fixed mask training are sequentially employed. CP and RIA can significantly improve the accuracy of V:N:M-sparse Transformers upon pruning, while for low training budget constraints, fixed mask training, no matter with full-parameters or LoRA, incurs significantly lower overhead compared to dynamic mask training as the mask update costs are canceled.
@@ 4.1 V and M selection
It is straightforward to prove that for the same Transformer, the relative order of different $MD_{f}$ is entirely determined by the values of V and M, i.e., $K$ in Eq. 5, and is irrespective of the weight shapes of the linear layers (See Appendix B for the proof). Thus, only calculating $K$ suffices for comparison and choosing the best (V, M) combination.
@@ 4.3 Three-staged LoRA training
where $\bm{\mathbf{B}}$ and $\bm{\mathbf{A}}$ are two low rank matrices. Normally, $\bm{\mathbf{M}}$ is a function of $\bm{\mathbf{W}}$, $\bm{\mathbf{B}}$, and $\bm{\mathbf{A}}$. During training, $\bm{\mathbf{W}}$ remains fixed while $\bm{\mathbf{B}}$ and $\bm{\mathbf{A}}$ are updated. We propose a three-stage LoRA training technique to enable dynamic mask training with LoRA and enhance the accuracy of V:N:M-sparse Transformers. 1) Dense LoRA. At the beginning of training, standard LoRA fine-tuning is applied to the Transformer, where the masks $\bm{\mathbf{M}}$ are consistently all-one matrices.
@@
2) Sparse LoRA with the dynamic masks. After the dense LoRA, the masks $\bm{\mathbf{M}}$ are updated at regular intervals according to the V:N:M sparse patterns as training progresses, which means each update occurs after a fixed number of iterations. During the mask update, the low-rank matrices $\bm{\mathbf{B}}$ and $\bm{\mathbf{A}}$ are merged with the original dense weight matrix $\bm{\mathbf{W}}$ before calculating the importance scores and conducting V:N:M-sparse pruning. 3) Sparse LoRA with the fixed masks. While dynamic mask updates facilitate the exploration of appropriate V:N:M-sparse masks, they can also introduce instability in the training process. Furthermore, directly applying regularization such as SR-STE, as illustrated in Eq. 3, is challenging because the LoRA term $\bm{\mathbf{B}}\bm{\mathbf{A}}$ is complex to regularize. Therefore, in the third stage, we advocate for fine-tuning with fixed masks to balance the exploration and exploitation of masks. At this stage, the masks $\bm{\mathbf{M}}$ are inherited from the last update in the previous stage and remain unchanged until the training is completed.
It is important to note that the number of iterations in the first two stages should constitute a smaller proportion of the total iterations, with the majority allocated to the third stage, i.e., Sparse LoRA with fixed masks. This is because, in the absence of regularization, frequent updates to the masks can negatively impact the gradient flow of V:N:M sparse Transformers during fine-tuning. As exemplified in Figure 3, the gradient norms of LoRA with dynmaic masks are consistently lower than those with fixed masks. In practice, the iterations for the first two stages should not exceed 10% of the total iterations. More details about our technique are shown in Appendix C .
@@ 5 Experiments
Models, datasets and tasks To evaluate the proposed V:N:M-sparse Transformer generation method—which incorporates V and M selection, V:N:M-specific channel permutations, and three staged LoRA training techniques, three benchmarks have been established: 1) DeiT (Touvron et al., 2021) for image classification. This benchmark is widely recognized for assessing the efficacy of model compression techniques in vision Transformers. Given that V:N:M sparsity operates at high sparsity levels (greater than 50%), the DeiT-tiny is excluded due to insufficient redundancy. The datasets used for the tasks include ImageNet-1K (Deng et al., 2009), Cifar-10 and Cifar-100 Krizhevsky et al. (2009), Bird and Vehicle from the subset of ImageNet-1K. Note that the latter four datasets are used to form downstream tasks. 2) Swin Transformers. This category of vision Transformers, known for its hierarchical architecture and shifted window mechanisms, demonstrates increased sensitivity to model compression (Liu et al., 2021). In this work, the V:N:M-sparse Swin Transformers are assessed across two tasks: image classification on the ImageNet-1K dataset and object detection on the COCO 2017 dataset (Lin et al., 2014). 3) Llama2-7B on downstream tasks including predicting the next token on wikitext2 (Merity et al., 2016) and eight well-established 5-shot tasks (Gao et al., 2021). In addition, the speedups of these V:N:M-sparse Transformers compared to their dense counterparts are measured on RTX 3090 GPUs, which were also the speed-testing platform in the initial study.
@@ 5.1 Results for V and M selection
Figure 4 compares the accuracy and speedup for V:N:M-sparse Transformers, both with and without the proposed V and M selection technique. In this experiment, for practical speedups, V is confined to $[16,32,64,128]$. Sparse Transformers with varying M values and unified V=64 are selected to establish the speedup threshold $s$, as specified in Eq. 4, with 64 representing a central value in the V distribution. Using the thresholds, the V and M selection technique is applied to derive new (V, M) and further generate the V:N:M-sparse Transformers accordingly, as indicated by the yellow lines in Figure 4. Specifically, in Figure 4(a), the V:N:M-sparse DeiT-base models located on the Pareto front exhibit higher Top-1 accuracy compared to the baseline, achieving a maximum accuracy improvement of 2.6% under similar speedup conditions. For the Llama2-7B, shown in Figure 4(b), the maximal average score difference is 3.41. Besides, when appropriate values of V and M are utilized, the sparse DeiT-base maintains a nearly lossless accuracy of 81.59% while achieving a speedup of 2.02 relative to its dense counterpart. Similarly, the sparse Llama2-7B achieves a speedup of 1.65, with a score of 50.85 on 5-shot tasks. Furthermore, it is essential to emphasize that V:N:M-sparse vision Transformers exhibit substantially greater speedups compared to 2:4 counterparts. As depicted in Figure 4(a), the 128:2:6-sparse DeiT-base matches the accuracy of the 2:4-sparse DeiT-base while achieving a 1.71x speedup over its dense counterpart, in contrast to the 2:4-sparse version, which achieves only a 1.15x speedup. More speedup results of V:N:M-sparse Transformers are shown in Appendix D and E.
Figure 4: The speedup-accuracy curves of V:N:M-sparse Transformer. a) Top-1 accuracy of DeiT-base using TS2 with different V and M values. Accuracy of dense DeiT-base: 81.84%. b) Average scores of Llama2-7B using TS3 on 5-shot tasks with different V and M values. Average score of dense Llama2-7B: 61.99. Average score of 2:4-sparse Llama2-7B: 50.76. The speedup-accuracy curves using TS1 are shown in Figure 10 in Appendix F. More results for larger Transformers are shown in Figure 11, 12 and 13 in Appendix G, respectively
@@ 5.2 Results for V:N:M-specific channel permutation and TS1
Since the proposed V:N:M-specific CP is effective for low-budget training scenarios, we assess this technique on downstream tasks that allow for a limited number of training epochs to attain acceptable accuracy. The results, presented in Tables 3 and 4, were all obtained using TS1. For vision downstream tasks with a maximum of 30 training epochs, our technique enhances final accuracy by 0.71% for full-parameter training and 1.43% for LoRA training. Notably, when V:N:M-sparse pruning is only applied without any training, the accuracy gap can increase to 5.87% with or without our technique. For Llama2-7B, which underwent only 1500 training iterations—a relatively short duration compared to the 50,000 iterations shown in the subsequent Table 8, our technique improves the 5-shot scores by 1.25. More results are shown in Appendix I.
@@ 5.3 Results for TS2
Table 5: Results comparison of sparse DeiTs on ImageNet-1K. See Appendix J for detailed description of (↓%), $\Delta$, *, and related works
. Model Param. (M) (↓%) FLOPs (G) (↓%) Top-1 Acc. (%) $\Delta$ DeiT-B 86.6 - 17.6 - 81.84 - DeiT-B-600e 86.6 - 17.6 - 82.01 +0.17 T2T-ViT-24 64.1 26.0 13.8 21.6 82.30 +0.46 PVT-L 61.4 29.1 9.8 44.3 81.70 -0.14 AutoFormer-B 54.0 37.6 11.0 37.5 82.40 +0.56 SSP-B 56.8 34.4 11.8 33.1 80.80 -1.04 S2ViTE-B 56.8 34.4 11.8 33.1 82.22 +0.38 Evo-ViT 56.3 - 11.7 33.3 80.30 -0.54 EViT-DeiT-B 86.6 - 11.5 34.7 81.50 -0.34 DynamicViT 61.0 - 11.5 34.7 81.30 -0.54 ViT-Slim 52.6 39.3 10.6 39.6 82.40 +0.56 VTP-B 47.3 45.4 10.0 43.2 80.70 -1.14 PS-ViT-B 86.6 - 9.8 44.3 81.50 -0.34 UVC - - 8.0 54.5 80.57 -1.27 SAViT 25.4 70.7 5.3 69.9 81.66 -0.18 NViT-B (ASP) 17 80.4 6.8 61.4 83.29 -0.07* Ours (64:2:8) 22.7 73.8 5.0 71.6 81.08 -0.76 Ours-600e (64:2:8) 22.7 73.8 5.0 71.6 81.76 -0.08 DeiT-S 22.1 - 4.6 - 79.85 - DeiT-S-600e 22.1 - 4.6 - 80.02 +0.17 AdaViT-S 22.1 0.0 3.6 21.7 78.60 -1.25 DynamicViT 22.1 - 3.4 26.1 79.60 -0.25 EViT-DeiT-S 22.1 - 3.4 34.8 79.50 -0.35 SSP-S 14.6 33.3 3.1 31.6 77.74 -2.11 S2ViTE-S 14.6 33.3 3.1 31.6 79.22 -0.63 SAViT 14.7 33.5 3.1 31.7 80.11 +0.26 NViT-S (ASP) 10.5 52.5 4.2 8.7 82.19 +0.99* Ours (64:2:5) 9.3 57.9 2.1 54.3 78.97 -0.68 Ours-600e (64:2:5) 9.3 57.9 2.1 54.3 79.65 -0.2
@@
Table 6: V:N:M-sparse Swin Transformers on image classification and object detection. See Appendix J for detailed description of (↓%), $\Delta$ and related works
. Model Method Param.(M) (↓%) FLOPs (G) (↓%) Top-1 Acuu. (%) Swin-base Dense 87.8 - 15.4 - 83.51 LPViT 64.1* 27.0 11.24 27.0 81.7 Ours(32:2:5) 35.9 59.1 6.1 60.4 81.7 H-DETR (Swin-Tiny) Dense 40.2 - 212.0 - 45.3 Dense-24e 40.2 - 212.0 - 49.2 FMT(32:2:5) 18.1 55.0 93.7 55.8 47.5 Ours(32:2:5) 18.1 55.0 93.7 55.8 47.8 FMT(32:2:6) 15.5 61.4 78.0 63.2 44.1 Ours(32:2:6) 15.5 61.4 78.0 63.2 45.3
@@ 5.4 Results for three-staged LoRA training and TS3
Figure 5: Average scores of V:N:M-sparse Llama2-7B on 5-shot tasks under different LoRA finetuning schemes. Dense: 61.99. 2:4-sparse: 50.76
@@ 6 Conclusion
This study focuses on enhancing the accuracy and accuracy-speedup trade-offs of V:N:M-sparse Transformers in multiple scenarios. We address the crucial yet unexplored questions specific to V:N:M sparsity, including selecting appropriate values for V and M, and CP tailored for V:N:M sparsity. Additionally, we propose a three-staged LoRA training technique, which for the first extends V:N:M sparsity to LLMs. Extensive experiments demonstrate that, with our methodology, V:N:M-sparse Transformers can attain nearly lossless accuracy or perform comparably to those with post-pruning 2:4 sparsity. Given its superior speed performance, we conclude that V:N:M sparsity is more effective than 2:4 for compressing highly redundant Transformers in inference-cost-sensitive scenarios. We hope our work to promote the widespread use of V:N:M sparsity as a truly effective solution for compressing Transformers.